OpenStudy (keke):

simplify; 8-b over 2b-16

OpenStudy (anonymous):

\[\frac{8-b}{-2*(8-b)}=-\frac{1}{2}\]

jhonyy9 (jhonyy9):

you want writing (8-b)/(2b-16) if yes this will be (8-b)/-2(8-b) --- 1/(-2) --- (-1/2)

OpenStudy (anonymous):

b = 8

OpenStudy (anonymous):

no no sorry

OpenStudy (keke):

simplify doesnt mean solve .

OpenStudy (anonymous):

i know keke that is y i said no no sorry understood first see what i wrote and then type